Obtain the Fourier series for the following function


f(x)=0 when -π<x<0


f(x)=sinx when 0<x<π


1
Expert's answer
2021-05-25T05:40:58-0400

First we calculate the constant 


"a_0=\dfrac{1}{\pi}\displaystyle\int_{-\pi}^{\pi}f(x)dx=\dfrac{1}{\pi}\displaystyle\int_{0}^{\pi} \sin xdx"

"=\dfrac{1}{\pi}\big[-\cos x\big]\begin{matrix}
 \pi \\
 0
\end{matrix}=-\dfrac{1}{\pi}(\cos \pi-\cos 0)=\dfrac{2}{\pi}"

Find the Fourier coefficients for "n\not=0"


"a_n=\dfrac{1}{\pi}\displaystyle\int_{-\pi}^{\pi}f(x)\cos nxdx"

"=\dfrac{1}{\pi}\displaystyle\int_{0}^{\pi} \sin x\cos(nx)dx"

"=\dfrac{1}{2\pi}\displaystyle\int_{0}^{\pi} (\sin( (n+1)x)-\sin((n-1)x))dx"

"=\dfrac{1}{2\pi}\big[-\dfrac{\cos( (n+1)x)}{n+1}+\dfrac{\cos( (n-1)x)}{n-1}\big]\begin{matrix}
 \pi \\
 0
\end{matrix}"

"=\dfrac{1}{2\pi}\big((-1)^n\dfrac{1}{n+1}-(-1)^n\dfrac{1}{n-1}+\dfrac{1}{n-1}-\dfrac{1}{n+1}\big)"

"=\dfrac{1}{2\pi}\big(-\dfrac{2}{2k+1+1}+\dfrac{2}{2k+1-1}\big)"

"=\dfrac{1}{2\pi}\big(\dfrac{1}{k}-\dfrac{1}{k+1}\big)"

"=\dfrac{1}{2\pi k(k+1)}"




"b_n=\dfrac{1}{\pi}\displaystyle\int_{-\pi}^{\pi}f(x)\sin nxdx"

"=\dfrac{1}{\pi}\displaystyle\int_{0}^{\pi} \sin x\sin(nx)dx"

"=\dfrac{1}{2\pi}\displaystyle\int_{0}^{\pi} (\cos( (n-1)x)-\cos((n+1)x))dx"

"=\dfrac{1}{2\pi}\big[\dfrac{\sin( (n-1)x)}{n-1}-\dfrac{\sin( (n+1)x)}{n+1}\big]\begin{matrix}
 \pi \\
 0
\end{matrix}"

"=\dfrac{1}{2\pi}\big(0-0-0+0\big)=0"


Thus, the Fourier series is


"f(x)=\dfrac{1}{\pi}+\dfrac{1}{2\pi k(k+1)}\cos((2k+1)x)"